[Ncep.nhc.nco_contacts] TSB Morning Rounds - Wed Jul 06, 2016

TSB Admin nhc.tsbadmin at noaa.gov
Wed Jul 6 15:49:45 UTC 2016


Here is today's summary of NHC computer operations:
--- Craig Mattocks

----------

TAFB:
1. N-AWIPS/AWIPS-2: Update on METEOSAT pseudo-color satellite products not  
ingesting. NHC receives the GEOColor satellite imagery from CIRA, which has  
been dealing with server problems. There was also a NESDIS/STAR outage for  
Himawari data starting at 00 UTC yesterday, which was resolved at 14 UTC.

2. GFE/NCP: Update on the limited availability of Altika/Cryosat  
(altimetry) data. Senior Hurricane Specialist Mike Brennan discovered that  
there was a data outage on the NESDIS/STAR ftp server, from which NHC pulls  
the data.

3. ATCF: TAFB forecasters have noticed glitches and missing data when  
launching the "safe mode" version of the ATCF on grunt, which runs on  
atcfsvr2. Sometimes launching the ATCF causes corruption in the display of  
other applications (eg, NMAP). These problems seem to be  
workstation-specific; they cannot be replicated on cumulus. They might be  
caused by a lack of video memory available or too many graphics-intensive  
applications running on a particular workstation. TSB Team ATCF is  
investigating.
HSU:
1. ATCF: Update on sluggishness of the Pacific desk ATCF workstation. NCO  
Linux Sysadmins Brian Maher and Shawn Gindhart have corrected a problem  
with a missing ~/.Xauthority file and will back the kernel down from  
release 642 to 573. Eventually the workstation will be upgraded from the  
RHEL 6.7 to RHEL 6.8 operating system.

2. ATCF: Update on incorrect timezones on ATCF workstations. NCO Linux  
Sysadmins Brian Maher and Shawn Gindhart have changed the timezone to UTC  
on the Pacific desk ATCF workstation, but it remains set to EDT at the  
Atlantic desk. The ~/.Xauthority file is still missing from the atcf-ops  
account on the Atlantic workstation.

3. ATCF: Update on delayed import of Dvorak fixes. ATCF dataflow expert  
Monica Bozeman will parallelize the fix staging script so it processes  
multiple basins simultaneously, instead of waiting for each basin to be  
processed in sequential fashion. This should result in more timely fix  
imports.

4. LDM: The TWO text products are not updating regularly in NWX. (For  
example, the latest TWOAT is from 18Z Tuesday, but the TWOEP is up to  
date.) TSB Developers Matt Sardi and Dave Zelinsky will check the LDM feed  
into $TEXT_DATA and $TEXT_DATA2.
Other items:
1. NOAA Environmental Satellite Processing Center (ESPC) firewall partner  
zone network maintenance is scheduled from 1400 to 1800 Z today. Incoming  
and outgoing data for ESPC products for NCEP, NAVO, FNMOC and AFWA data  
connections will be impacted. Only brief periods of intermittent  
connectivity loss is expected.
NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160706/6b08fd12/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list